Ruthenium-106 has a half-life of 1.0 years. How many half-lives will pass in 25 years?


1
Expert's answer
2021-06-09T11:56:02-0400

At very first assume that the solution is of "1l"

Now:

Comparing the "K_a" of acetic acid, "1.76X10^-5" , these solutions aren’t very large comparative to .25M and .1M, it’s a buffer equation because there is a acid and conjugate base, so you’re able to use the Henderson Hasselbalch equation.


Now :


"P_H= P_{Ka} + log (\dfrac{base}{acid})"


"P_H= 4.76 + log(\dfrac{0.1}{0.25})"


"P_H=4.362"